OpenStudy (anonymous):

If a1=7 a4=26 for arithmetic series find the missing term . Repeat for. Geometric series HELP PlEASE :(

OpenStudy (dumbcow):

a4 = a1 + d(4-1) 26 = 7 +3d d = 19/3 a_n = a1+ (19/3)(n-1)

OpenStudy (dumbcow):

a4 = a1*r^4-1 26 = 7*r^3 r = cube root of 26/7 a_n = 7*(26/7)^(n-1)/3

OpenStudy (anonymous):

Wow thanks <33